### v2

# Branched covers between $2$-spheres

Abstract.
By leveraging techniques from complex analysis and topology, we have uncovered a remarkable property of branched covers and formulated a criterion for the existence of specific types of branched covers between 2-spheres. This criterion extends the findings previously reported by Jiang (2004), Pervova-Petronio (2006), Zhu (2019), and Wei-Wu-Xu (2024). Furthermore, as an application of our criterion, we introduce several novel exceptional branching datum.
Key words   Branched cover, branching datum, the Hurwitz existence problem.
2020 MR Subject Classification:   57M12.
1. Introduction
Investigating the existence of branched covers between two compact, connected and orientable surfaces, given a specific branching datum, is a famous problem within the field of complex analysis and topology. It dates back to Hurwitz [6] and is known as the Hurwitz existence problem. This problem exemplifies the interconnectedness of mathematics, as it intertwines algebra, geometry, topology, and analysis, remaining a topic of ongoing research.
Considerable energy has been devoted over the time to attack the Hurwitz existence problem. In 1891, Hurwitz in his celebrated paper [6] transformed the existence problem of branched covers into the existence problem of certain permutations related to the symmetric group and partitions. In 1984, Edmonds, Kulkarni and Stong [5] proved that all candidate branching datum were realizable when the target space $X$ satisfies $\chi(X)\leq 0$. However, When the target space $X$ satisfies $\chi(X)=2$, i.e. $X=S^{2}$ the Riemann sphere, the problem becomes very complex, and there exist exceptional cases. For example, this is the case for a abstract branching datum $\mathcal{D}=\{[3,1],[2,2],[2,2]\}$. In [5], Edmonds, Kulkarni and Stong proposed the prime degree conjecture and reduced this conjecture to a scenario with only three branching points. Pakovich [10] and Pascali-Petronio [12, 13] provided strong evidence in support of the conjecture, respectively. However, the conjecture remains open to this day. In 2001, Barański [1] employed graph theory on surfaces to determine which branching datum could be realized through branched covers of the Riemann sphere $S^{2}$. In 2004, Jiang [7] used differential topology to provide a sufficient and necessary condition for the realization of some classes of abstract branching datum. In 2006, Zheng [24] determined all exceptional branching datum by computer when the target space is the Riemann sphere $S^{2}$ and the source spaces are the Riemann sphere $S^{2}$ or torus $T^{2}$ with three branching points and degree $d\leq 22$. In the same year, Pervova and Petronio [14] provided examples of partially realizable branching datum and exceptional branching datum based on three different techniques: dessins d’enfants, decomposability, and graphs on surfaces. Recently, Pakovich [11] constructed a large number of new such branching datum using results and techniques related to fiber products of holomorphic maps between compact Riemann surfaces. For further results, we refer the interested reader to [2, 5, 20, 17] and the references cited therein. Particularly, we refer the reader to [3, 15, 16] for a review of available results and techniques.
Describing the branching datum of all branched covers remains an extremely complex and elusive task. In 2024, Wu, Xu, and the second author [21] established a property of branched covers on the Riemann sphere with three branching points and introduced several new examples of exceptional branching datum. In this paper, we continue to investigate the existence of branched covers between Riemann spheres with at least three branching points. By leveraging Stoïlow’s theorem and a well known result from differential topology, we discover a remarkable property (Theorem 1.1) of a specific type of branched cover. Then we transform the problem of the existence of branched covers into the problem of the existence of rational maps between Riemann spheres. Consequently, we have derived a criterion for the existence of branched covers between Riemann spheres that have at least three branching points. Our findings are summarized as follows:
Theorem 1.1.
Suppose $s\geq 2,t\geq 1$ and $d^{\prime}\geq 1$ are integers. Suppose $A_{1}=[sa_{11},sa_{12},\ldots,sa_{1r_{1}}],A_{2}=[sa_{21},sa_{22},\ldots,sa_{2r_{2}}],A_{3}=[ta_{31},ta_{32},\ldots,ta_{3r_{3}}]$, $\cdots,A_{n}=[a_{n1},a_{n2},\ldots,a_{nr_{n}}]$ are $n\geq 3$ nontrivial partitions of $sd^{\prime}$, where all $a_{kl}\geq 1$ are integers. If there exists a branched cover $f:S^{2}\rightarrow S^{2}$ between $2$-spheres such that its branching datum is
$\mathcal{D}=\{A_{1},A_{2},A_{3},\ldots,A_{n}\},$
then the following cases hold:
  1. (1)
    When $s\geq 4$, then $t=1$.
  2. (2)
    When $s=3$, then $t$ is either $1$ or $2$. Furthermore, when $t=2$, then $4|d^{\prime}$.
  3. (3)
    When $s=2$, then $t\geq 1$. Furthermore, when $t\geq 2$, then $t|d^{\prime}$.
Theorem 1.2.
Let $s\geq 2$ and $d^{\prime}\geq 1$ be integers. Let $A_{1}=[sa_{11},\ldots,sa_{1r_{1}}]$, $A_{2}=[sa_{21},\ldots,sa_{2r_{2}}],A_{3}=[a_{31},\ldots,a_{3r_{3}}],\ldots,A_{n}=[a_{n1},\ldots,a_{nr_{n}}]$ be $n\geq 3$ nontrivial partitions of $sd^{\prime}$ with the condition that $\sum\limits_{k=1}^{n}r_{k}=s(n-2)d^{\prime}+2$. Then, there exists a branched cover $f:S^{2}\rightarrow S^{2}$ between $2$-spheres with degree $sd^{\prime}$ and branching datum
$\mathcal{D}=\{A_{1},A_{2},A_{3},\ldots,A_{n}\},$
if and only if there exists a branched cover $g:S^{2}\rightarrow S^{2}$ with degree $d^{\prime}$ such that its branching datum consists of $[a_{11},\ldots,a_{1r_{1}}]$ and $[a_{21},\ldots,a_{2r_{2}}]$, together with data arising by splitting the partition of $A_{k},k\geq 3$ into $s$ partitions of $d^{\prime}$.
Theorem 1.3.
Let $d^{\prime}\geq 2$ and $t\geq 2$ be integers such that $t|d^{\prime}$. Let $A_{1}=[2a_{11},\ldots,2a_{1r_{1}}]$, $A_{2}=[2a_{21},\ldots,2a_{2r_{2}}],A_{3}=[ta_{31},\ldots,ta_{3r_{3}}]$, $\ldots,A_{n}=[a_{n1},\ldots,a_{nr_{n}}]$ be $n\geq 3$ nontrivial partitions of $2d^{\prime}$ with the condition that $\sum\limits_{k=1}^{n}r_{k}=2(n-2)d^{\prime}+2$. Then, there exists a branched cover $f:S^{2}\rightarrow S^{2}$ between $2$-spheres with degree $2d^{\prime}$ and branching datum
$\mathcal{D}=\{A_{1},A_{2},A_{3},\ldots,A_{n}\},$
if and only if there exists a branched cover $g:S^{2}\rightarrow S^{2}$ with degree $\frac{d^{\prime}}{t}$ such that its branching datum consists of partitions $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$ and $[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, together with data arising by splitting the partition of $[a_{11},\ldots,a_{1r_{1}}]$ and $[a_{21},\ldots,a_{2r_{2}}]$ into $t$ partitions of $\frac{d^{\prime}}{t}$, and splitting the partition of $A_{k},k\geq 4$ into $2t$ partitions of $\frac{d^{\prime}}{t}$. Here $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$ and $[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$ are two partitions of $\frac{d^{\prime}}{t}$ obtained by splitting the partition of $[a_{31},\cdots,a_{3r_{3}}]$.
Theorem 1.4.
Let $d^{\prime}\geq 4$ be an integer with $4|d^{\prime}$. Let $A_{1}=[3a_{11},\ldots,3a_{1r_{1}}]$, $A_{2}=[3a_{21},\ldots,3a_{2r_{2}}],A_{3}=[2a_{31},\ldots,2a_{3r_{3}}]$, $\ldots,A_{n}=[a_{n1},\ldots,a_{nr_{n}}]$ be $n\geq 3$ nontrivial partitions of $3d^{\prime}$ with the condition that $\sum\limits_{k=1}^{n}r_{k}=3(n-2)d^{\prime}+2$. Then, there exists a branched cover $f:S^{2}\rightarrow S^{2}$ between $2$-spheres with degree $3d^{\prime}$ and branching datum
$\mathcal{D}=\{A_{1},A_{2},A_{3},\ldots,A_{n}\},$
if and only if there exists a branched cover $g:S^{2}\rightarrow S^{2}$ with degree $\frac{d^{\prime}}{4}$ such that its branching datum consists of the partitions $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}],[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, $[\widetilde{a}^{3}_{31},\ldots,\widetilde{a}^{3}_{3\widetilde{r}_{3}}]$, $[\widetilde{a}^{4}_{31},\ldots,\widetilde{a}^{4}_{3\widetilde{r}_{4}}]$, $[\widetilde{a}^{5}_{31},\ldots,\widetilde{a}^{5}_{3\widetilde{r}_{5}}]$ and $[\widetilde{a}^{6}_{31},\ldots,\widetilde{a}^{6}_{3\widetilde{r}_{6}}]$, together with data arising by splitting the partition $[a_{11},a_{12},\ldots,a_{1r_{1}}]$ and $[a_{21},a_{22},\ldots,a_{2r_{2}}]$ into $4$ partitions of $\frac{d^{\prime}}{4}$, and splitting the partition of $A_{k},k\geq 4$ into $12$ partitions of $\frac{d^{\prime}}{4}$. Here $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}],[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, $[\widetilde{a}^{3}_{31},\ldots,\widetilde{a}^{3}_{3\widetilde{r}_{3}}]$, $[\widetilde{a}^{4}_{31},\ldots,\widetilde{a}^{4}_{3\widetilde{r}_{4}}]$, $[\widetilde{a}^{5}_{31},\ldots,\widetilde{a}^{5}_{3\widetilde{r}_{5}}]$ and $[\widetilde{a}^{6}_{31},\ldots,\widetilde{a}^{6}_{3\widetilde{r}_{6}}]$ are $6$ partitions of $\frac{d^{\prime}}{4}$ obtained by splitting the partition of $[a_{31},\cdots,a_{3r_{3}}]$.
Remark 1.1.
When $n=3$, the statement in Theorem 1.2 specializes to the result of Wei-Wu-Xu [21]. When $s=2$ and $n\geq 3$, the statement in Theorem 1.2 specializes a result of Perova-Petronio [14]. Furthermore, Theorem 1.3 is equivalent to the result of Jiang [7].
By these results, one would have immediate corollaries for realizable data between $2$-spheres.
Corollary 1.1.
Under the conditions of theorem 1.2, the subsequent statements are valid.
  1. (1)
    Each $\alpha_{j}^{i}\leq d^{\prime}$.
  2. (2)
    If $r_{i}=2$ for some $i\geq 3$, then $s=2$ and $A_{i}=[d^{\prime},d^{\prime}]$.
  3. (3)
    If $r_{i}=s$ for some $i\geq 3$, then $A_{i}=[\underbrace{d^{\prime},\ldots,d^{\prime}}_{s}]$.
Corollary 1.2.
Under the conditions of theorem 1.3, the subsequent statements are valid.
  1. (1)
    Each $\alpha_{j}^{i}\leq d^{\prime}$.
  2. (2)
    If $r_{i}=4$ for some $i\geq 4$, then $t=2$ and $A_{i}=[\frac{d^{\prime}}{2},\frac{d^{\prime}}{2},\frac{d^{\prime}}{2},\frac{d^{\prime}}{2}]$.
  3. (3)
    If $r_{i}=2t$ for some $i\geq 4$, then $A_{i}=[\underbrace{\frac{d^{\prime}}{t},\ldots,\frac{d^{\prime}}{t}}_{2t}]$.
Corollary 1.3.
Under the conditions of theorem 1.4, the subsequent statements are valid.
  1. (1)
    Each $\alpha_{j}^{i}\leq d^{\prime}$.
  2. (2)
    If $r_{i}=6$ for some $i\geq 4$, then $A_{i}=[\frac{d^{\prime}}{2},\frac{d^{\prime}}{2},\frac{d^{\prime}}{2},\frac{d^{\prime}}{2},\frac{d^{\prime}}{2},\frac{d^{\prime}}{2}]$.
Remark 1.2.
A natural question arises: Suppose there exists a branched cover $f:S^{2}\rightarrow S^{2}$ with branching datum
$\mathcal{D}=\{[\alpha_{1},\cdots,\alpha_{A}],[\beta_{1},\cdots,\beta_{B}],[\gamma_{1},\cdots,\gamma_{C}]\}.$
We wonder whether, for any candidate branching datum
$\widetilde{\mathcal{D}}=\{[\alpha_{1},\cdots,\alpha_{A}],[\beta_{1},\cdots,\beta_{B}],[\widetilde{\gamma}_{1},\cdots,\widetilde{\gamma}_{C}]\}$
there exists a branched cover $g:S^{2}\rightarrow S^{2}$ such that its branching datum is $\widetilde{\mathcal{D}}$. The answer is negative. For instance, as shown in [24], there exists a branched cover with the branching datum
$\{[5,3],[2,2,2,2],[3,2,2,1]\},$
but there does not exist a branched cover with the branching datum
$\{[5,3],[2,2,2,2],[3,3,1,1]\}.$
Finally, by utilizing the result of Song-Xu presented in [17] (Theorem 1.3), one can readily demonstrate the following theorem on existence.
Theorem 1.5.
Suppose $k\geq 3,x\geq 1$ and $y\geq 1$ are integers with $k\geq x$ and $k\geq y$. There exists a rational map $f:\overline{\mathbb{C}}\rightarrow\overline{\mathbb{C}}$ between Riemann sphere with branching datum
$\{[a_{1},a_{2},\ldots,a_{x+y}],[\underbrace{2,\ldots,2}_{k-y},2y],[\underbrace{2,\ldots,2}_{k-x},2x]\}$
if and only if $[a_{1},a_{2},\ldots,a_{x+y}]$ can be split into two partition of $k$ and $\frac{k}{{\rm GCD}(a_{1},a_{2},\ldots,a_{x+y})}\geq\max\{x,y\}$
The paper is structured as follows: Section 2 introduces some key concepts and necessary results. Section 3 provides the proofs of Theorems 1.2, 1.3 and 1.4, respectively. Section 4 presents some new examples of exceptional branching datum.
2. Preliminaries
In this section, we establish some necessary notations for branched covers between two compact, connected and orientable surfaces and present some necessary results that will be crucial for proving our findings. These notations and results can be referenced in the paper [12].
2.1. Branched covers
Let $Y$ and $X$ be two compact, connected and orientable surfaces. A branched cover is a map $f:Y\rightarrow X$ that is locally modelled on maps of the form $\mathbb{C}\ni z\mapsto z^{k}\in\mathbb{C}$ for some positive integer $k$. The integer $k$ is referred to as the local degree at the point of $Y$ corresponding to 00 in the source $\mathbb{C}$. If $k>1$ then the point of $X$ corresponding to 00 in the target $\mathbb{C}$ is termed a branching point. Branching points are isolated, implying there are finitely many of them, say $n$, of them. By removing the branching points in $X$ and all their pre-images in $Y$, the restriction of $f$ yields a genuine cover, the degree if which we denote by $d$. The collection of local degrees at the preimages of the $i$-th branching point of $X$ forms a nontrivial partition $A_{i}$ of $d$, which is a collection of positive integers summing up to $d$ with at least one element greater than $1$. We define $l(A_{i})$ as the length of this partition and $\mathcal{D}(f)=\{A_{1},A_{2},\ldots,A_{n}\}$ as the branching datum of $f$.
2.2. The Hurwitz existence problem
Let $f:Y\rightarrow X$ be a branched cover between two compact, connected and orientable surfaces $Y$ and $X$. Denote the Euler characteristics of $X$ and $Y$ by $\chi(X)$ and $\chi(Y)$, respectively. The degree $d$ of $f$, the number $n$ of branching points of $f$, and the Euler characteristics $\chi(X)$ and $\chi(Y)$ satisfy the famous Riemann-Hurwitz formula:
(1) $nd-\sum\limits_{i=1}^{n}l(A_{i})=d\chi(X)-\chi(Y).$
The Hurwitz existence problem asks whether, given two compact, connected and orientable surfaces $Y$ and $X$ with Euler characteristics $\chi(X)$ and $\chi(Y)$ respectively, an integer $d\geq 2$, and $n$ nontrivial partitions $A_{1},\ldots,A_{n}$ of $d$ which satisfy the Riemann-Hurwitz formula, there exists a branched cover $f:Y\rightarrow X$ such that $\mathcal{D}=\{A_{1},\ldots,A_{n}\}$ is the branching datum of $f$.
We usually refer to a set $\mathcal{D}$ of partitions of an integer $d$ as an abstract branching datum. A branching datum $\mathcal{D}$ that fulfills the Riemann-Hurwitz formula is called a candidate branching datum. If the solution to the Hurwitz existence problem is affirmative, we call $\mathcal{D}$ realizable; otherwise, it is termed exceptional.
2.3. The Riemann existence theorem
In 1891, Hurwitz [6] established the renowned the Riemann existence theorem, which transformed the existence problem of branched covers into the existence problem of certain permutations related to the symmetric group and partitions.
Theorem 2.1 ([6]).
Let $X$ be a connected Riemann surface and $\Delta$ a discrete subset of $X$. Given $d\geq 1$ and a transitive permutation representation $\rho:\pi_{1}(X\backslash\Delta)\rightarrow S_{\{1,2,\ldots,d\}}$, the symmetry group of $\{1,2,\ldots,d\}$, there exists a connected Riemann surface $Y$ and a proper holomorphic map $f:Y\rightarrow X$ which realises $\rho$ as its monodromy homomorphism. Moreover, $Y$ and $f$ are unique up to equivalence.
2.4. Stoïlow’s theorem
Recall that a continuous map $f:Y\rightarrow X$ between topological spaces is light if the pre-image $f^{-1}(x)$ of each point $x\in X$ is totally disconnected, and discrete if $f^{-1}(x)$ is a discrete subset of $Y$. A continuous map is open if the image of each open set is an open set. The branch set $B_{f}$ of a continuous map $f:Y\rightarrow X$ is the set of points $y\in Y$ at which $f$ fails of be a local homeomorphism. Stoïlow’s theorem from 1928 states that a continuous, open, and light map between surfaces is a discrete map with a discrete branch set. This result implies that such maps between orientable surfaces are locally modeled by power maps $z\rightarrow z^{k}$ and admit a holomorphic factorization. This local description indicates a deep connection between continuous, open and light maps and holomorphic maps between surfaces. This connection was found by Stoïlow [19] and Whyburn [22, 23]:
Theorem 2.2 ([8]).
For a continuous, open and light map $f:Y\rightarrow X$ between Riemann surfaces there exists a Riemann surface $\widetilde{Y}$ and a homeomorphism $h:Y\rightarrow\widetilde{Y}$ for which $f\circ h^{-1}:\widetilde{Y}\rightarrow X$ is a holomorphic map.
3. proofs of theorems
3.1. Three lemmas
In this section, we will establish several lemmas that will be repeatedly utilized in the proofs of the theorems. First, we recall a well known result in the field of topology.
Theorem 3.1.
Let $X$ be a connected smooth manifold with dimension $dimX\geq 2$. Let $p_{1},\cdots,p_{n}$ be $n$ distinct points in $X$. Then for any $n$ distinct points $q_{1},\cdots,q_{n}$ in $X$ there exists a smooth homeomorphism $h:X\rightarrow X$ that is homotopic to the identity map and satisfies $h(p_{i})=q_{i}$, for $i=1,\cdots,n$.
From Theorems 3.1 and 2.2, we obtain the following lemmas:
Lemma 3.1.
Let $f:Y\rightarrow X$ be a branched cover between two compact connected Riemann surfaces $Y$ and $X$ with branching datum $\mathcal{D}=\{A_{1},\cdots,A_{n}\}$. Suppose the branching points of $f$ in $X$ are $p_{1},\cdots,p_{n}$. Then for any $n$ distinct points $q_{1},\cdots,q_{n}$ in $X$ there exists a branched cover $g:Y\rightarrow X$ such that its branching datum is $\mathcal{D}=\{A_{1},\cdots,A_{n}\}$ and its branching points are $q_{1},\cdots,q_{n}$.
Proof.
By Theorem 3.1, there exists a smooth homeomorphism $h:X\rightarrow X$ such that $h(p_{i})=q_{i}$ for $i=1,\cdots,n$. Obviously, $h\circ f:Y\rightarrow X$ is the desired map. ∎
Lemma 3.2.
Let $f:Y\rightarrow X$ be a branched cover between two compact connected Riemann surfaces $Y$ and $X$. Then there exists a Riemann surface $\widetilde{Y}$ and a homeomorphism $h:Y\rightarrow\widetilde{Y}$ such that the composition $f\circ h^{-1}:\widetilde{Y}\rightarrow X$ is a holomorphic map.
Proof.
This result is directly derived from Theorem 2.2. ∎
Furthermore, when $Y=X=S^{2}$, given that the complex structure on $S^{2}$ is unique, we can deduce the following lemma.
Lemma 3.3.
Let $f:S^{2}\rightarrow S^{2}$ be a branched cover between the Riemann spheres $S^{2}$ with $n$ branching points. Then for any distinct $n$ points $p_{1},\ldots,p_{n}$ in $S^{2}$ there exists a holomorphic branched cover $g:S^{2}\rightarrow S^{2}$ such that its branching datum is identical to that of $f$, and $p_{1},\ldots,p_{n}$ are branching points of $g$.
3.2. Proof of Theorem 1.1
Proof.
Given the existence of a branched cover $f:S^{2}\rightarrow S^{2}$ with degree $sd^{\prime}$ and branching datum $\mathcal{D}=\{A_{1},A_{2},A_{3},\ldots,A_{n}\}$, by the Riemann-Hurwitz formula, we deduce that
$\sum\limits_{k=1}^{n}r_{k}=s(n-2)d^{\prime}+2.$
Noting the inequalities
$r_{1}\leq d^{\prime},~{}~{}r_{2}\leq d^{\prime},~{}~{}r_{3}\leq\frac{sd^{\prime}}{t},~{}~{}r_{4}\leq sd^{\prime}-1,~{}~{}\cdots,~{}~{}r_{n}\leq sd^{\prime}-1,$
we deduce that
$s(n-2)d^{\prime}+2=\sum\limits_{k=1}^{n}r_{k}\leq 2d^{\prime}+\frac{sd^{\prime}}{t}+(n-3)(sd^{\prime}-1),$
which simplifies to
$n-1\leq(2+\frac{s}{t}-s)d^{\prime}.$
Define
$\tau(s,t)=2+\frac{s}{t}-s,~{}~{}s\geq 2,~{}~{}t\geq 1.$
Given that $0<n-1\leq(2+\frac{s}{t}-s)d^{\prime}=\tau(s,t)d^{\prime}$, we deduce that $\tau(s,t)>0$. However, when $s\geq 3$ and $t\geq 3$, $\tau(s,t)\leq\tau(3,3)=0$, which contradicts the condition $\tau(s,t)>0$. Therefore, when $s\geq 3$, $t$ must be either $1$ or $2$. When $s\geq 4$ and $t=2$, $\tau(s,2)=2-\frac{s}{2}\leq 0$, which also contradicts the condition $\tau(s,t)>0$. Hence, when $s\geq 4$, $t$ must be $1$. When $s=2$, then for all $t\geq 1$, $\tau(2,t)=\frac{2}{t}>0$. Consequently, we have identified the following three cases:
  1. (1)
    When $s\geq 4$, then $t=1$.
  2. (2)
    When $s=3$, then $t$ is either $1$ or $2$.
  3. (3)
    When $s=2$, then $t\geq 1$.
When $s=3$ and $t=2$, we have
$A_{1}=[3a_{11},\ldots,3a_{1r_{1}}],A_{2}=[3a_{21},\ldots,3a_{2r_{2}}],A_{3}=[2a_{31},\ldots,2a_{3r_{3}}].$
Regard $S^{2}$ as the Riemann sphere ${}\mathbb{C}\cup\{\infty\}$. By Lemma 3.3, we can assume $f:S^{2}\rightarrow S^{2}$ is rational map and the branching points corresponding to the partitions $A_{1}$ and $A_{2}$ are $0,\infty$, respectively. Since $3$ is a common divisor of the elements of the partitions $A_{1}$ and $A_{2}$, we know that there exists a rational map $h:S^{2}\rightarrow S^{2}$ with degree $d^{\prime}$ such that its branching datum consists of partitions $[a_{11},\ldots,a_{1r_{1}}]$ and $[a_{21},\ldots,a_{2r_{2}}]$, together with data arising by splitting the partition $A_{k},k\geq 3$ into $3$ partitions of $d^{\prime}$.
Denote the $3$ partitions of $d^{\prime}$ obtained by splitting the partition $A_{3}$ by
$A_{31}=[2a^{1}_{31},\ldots,2a^{1}_{3r_{1}}],A_{32}=[2a^{2}_{31},\ldots,2a^{2}_{3r_{2}}],A_{33}=[2a^{3}_{31},\ldots,2a^{3}_{3r_{3}}].$
Since $2$ is a common divisor of the elements of the partitions $A_{31}$ and $A_{32}$, by Lemma 3.3, we know that there exists a rational map $g:S^{2}\rightarrow S^{2}$ with degree $\frac{d^{\prime}}{2}$ such that its branching datum consists of partitions $[a^{1}_{31},\ldots,a^{1}_{3r_{1}}]$ and $[a^{2}_{31},\ldots,a^{2}_{3r_{2}}]$, together with data arising by splitting the partition $A_{k},k\geq 4$ into $6$ partitions of $\frac{d^{\prime}}{2}$, and splitting the partition $A_{33}$ into $2$ partitions of $\frac{d^{\prime}}{2}$.
Denote the $2$ partitions of $\frac{d^{\prime}}{2}$ obtained by splitting the partition $A_{33}$ by
$A_{331}=[2\widetilde{a}^{1}_{31},\ldots,2\widetilde{a}^{1}_{3\widetilde{r}_{1}}],A_{332}=[2\widetilde{a}^{2}_{31},\ldots,2\widetilde{a}^{2}_{3\widetilde{r}_{2}}].$
Since $2$ is a common divisor of the elements of the partitions $A_{331}$ and $A_{332}$, by Lemma 3.3, we know that there exists a rational map of degree $\frac{d^{\prime}}{4}$. Consequently, we deduce that $4|d^{\prime}$.
Similarly, when $s=2$ and $t\geq 2$, we can deduce that $t|d^{\prime}$. ∎
3.3. Proof of Theorem 1.2
Proof.
The necessity is established in Theorem 1.1. Now, we proceed to prove the sufficiency.
Suppose there exists a branched cover $g:S^{2}\rightarrow S^{2}$ with degree $d^{\prime}$ such that its branching datum consists of the partitions $[a_{11},\ldots,a_{1r_{1}}]$ and $[a_{21},\ldots,a_{2r_{2}}]$, along with data arising from splitting the partition of $A_{k}$ for $k\geq 3$ into $s$ partitions of $d^{\prime}$.
Regard $S^{2}$ as the Riemann sphere ${}\mathbb{C}\cup\{\infty\}$. By Lemma 3.3, we can assume that $g$ is a rational map with branching points at 00 and $\infty$ corresponding the partitions $[a_{11},\ldots,a_{1r_{1}}]$ and $[a_{21},\ldots,a_{2r_{2}}]$, respectively, and the other branching points are assigned suitable complex numbers. By direct calculation, $g^{s}$ is the desired map. ∎
3.4. Proof of Theorem 1.3
Proof.
The necessity is established in Theorem 1.1. Now, we proceed to prove the sufficiency.
Suppose there exists a branched cover $g:S^{2}\rightarrow S^{2}$ with degree $\frac{d^{\prime}}{t}$ such that its branching datum consists of partitions $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$ and $[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, along with data arising by splitting the partition $[a_{11},\ldots,a_{1r_{1}}]$ and $[a_{21},\ldots,a_{2r_{2}}]$ into $t$ partitions of $\frac{d^{\prime}}{t}$, and splitting the partition of $A_{k}$ for $k\geq 4$ into $2t$ partitions of $\frac{d^{\prime}}{t}$. Here $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$ and $[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$ are two partitions of $\frac{d^{\prime}}{t}$ obtained by splitting the partition of $[a_{31},\cdots,a_{3r_{3}}]$.
Regard $S^{2}$ as the Riemann sphere ${}\mathbb{C}\cup\{\infty\}$. By Lemma 3.3, we can assume that $g$ is a rational map with branching points at 00 and $\infty$ corresponding the partitions $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$ and $[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, respectively, while the other branching points are assigned suitable complex numbers. Then $g^{t}$ is a rational map with branching datum $[a_{11},\ldots,a_{1r_{1}}]$, $[a_{21},\ldots,a_{2r_{2}}]$, $[t\widetilde{a}^{1}_{31},\ldots,t\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$, $[t\widetilde{a}^{2}_{31},\ldots,t\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, and the data arising from splitting the partition of $A_{k}$ for $k\geq 4$ into $2$ partitions of $d^{\prime}$. By Lemma 3.3, there exists a rational map $h:S^{2}\rightarrow S^{2}$ with branching points at $-1$ and $1$ corresponding the partitions $[t\widetilde{a}^{1}_{31},\ldots,t\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$ and $[t\widetilde{a}^{2}_{31},\ldots,t\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, and branching points at 00 and $\infty$ corresponding the partitions $[a_{11},\ldots,a_{1r_{1}}]$ and $[a_{21},\ldots,a_{2r_{2}}]$, respectively, while the other branching points are assigned suitable complex numbers. By direct calculation, $h^{2}$ is the desired map.
3.5. Proof of Theorem 1.4
Proof.
The necessity is established in Theorem 1.1. Now, we proceed to prove the sufficiency.
Suppose there exists a branched cover $g:S^{2}\rightarrow S^{2}$ with degree $\frac{d^{\prime}}{4}$ such that its branching datum consists of the partitions $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}],[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, $[\widetilde{a}^{3}_{31},\ldots,\widetilde{a}^{3}_{3\widetilde{r}_{3}}]$, $[\widetilde{a}^{4}_{31},\ldots,\widetilde{a}^{4}_{3\widetilde{r}_{4}}]$, $[\widetilde{a}^{5}_{31},\ldots,\widetilde{a}^{5}_{3\widetilde{r}_{5}}]$ and $[\widetilde{a}^{6}_{31},\ldots,\widetilde{a}^{6}_{3\widetilde{r}_{6}}]$, along with data arising from splitting the partitions $[a_{11},\ldots,a_{1r_{1}}]$ and $[a_{21},\ldots,a_{2r_{2}}]$ into $4$ partitions of $\frac{d^{\prime}}{4}$, and splitting the partition of $A_{k}$ for $k\geq 4$ into $12$ partitions of $\frac{d^{\prime}}{4}$. Here $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}],[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, $[\widetilde{a}^{3}_{31},\ldots,\widetilde{a}^{3}_{3\widetilde{r}_{3}}]$, $[\widetilde{a}^{4}_{31},\ldots,\widetilde{a}^{4}_{3\widetilde{r}_{4}}]$, $[\widetilde{a}^{5}_{31},\ldots,\widetilde{a}^{5}_{3\widetilde{r}_{5}}]$ and $[\widetilde{a}^{6}_{31},\ldots,\widetilde{a}^{6}_{3\widetilde{r}_{6}}]$ are $6$ partitions of $\frac{d^{\prime}}{4}$ obtained by splitting the partition of $[a_{31},\cdots,a_{3r_{3}}]$.
By Lemma 3.3, we can assume $g$ is a rational map with branching points at 00 and $\infty$ corresponding the partitions $[\widetilde{a}^{1}_{31},\ldots,\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$ and $[\widetilde{a}^{2}_{31},\ldots,\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, respectively, branching points at $-1$ and $1$ corresponding the partitions $[\widetilde{a}^{3}_{31},\ldots,\widetilde{a}^{3}_{3\widetilde{r}_{3}}]$ and $[\widetilde{a}^{4}_{31},\ldots,\widetilde{a}^{4}_{3\widetilde{r}_{4}}]$, respectively, branching points at $-\sqrt{-1}$ and $\sqrt{-1}$ corresponding the partitions $[\widetilde{a}^{5}_{31},\ldots,\widetilde{a}^{5}_{3\widetilde{r}_{5}}]$ and $[\widetilde{a}^{6}_{31},\ldots,\widetilde{a}^{6}_{3\widetilde{r}_{6}}]$, respectively, with other branching points assigned suitable complex numbers. Then $g^{2}$ is a rational map with branching datum $[2\widetilde{a}^{1}_{31},\ldots,2\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$, $[2\widetilde{a}^{2}_{31},\ldots,\\ 2\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, $[\widetilde{a}^{3}_{31},\ldots,\widetilde{a}^{3}_{3\widetilde{r}_{3}},\widetilde{a}^{4}_{31},\ldots,\widetilde{a}^{4}_{3\widetilde{r}_{4}}]$, and $[\widetilde{a}^{5}_{31},\ldots,\widetilde{a}^{5}_{3\widetilde{r}_{5}},\widetilde{a}^{6}_{31},\ldots,\widetilde{a}^{6}_{3\widetilde{r}_{6}}]$.
By Lemma 3.3, there exists a rational map $h:S^{2}\rightarrow S^{2}$ with branching points at $-1$ and $1$ corresponding the partitions $[2\widetilde{a}^{1}_{31},\ldots,2\widetilde{a}^{1}_{3\widetilde{r}_{1}}]$ and $[2\widetilde{a}^{2}_{31},\ldots,2\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, respectively, with branching points at 00 and $\infty$ corresponding the partitions $[\widetilde{a}^{3}_{31},\ldots,\widetilde{a}^{3}_{3\widetilde{r}_{3}},\widetilde{a}^{4}_{31},\ldots,\widetilde{a}^{4}_{3\widetilde{r}_{4}}]$ and $[\widetilde{a}^{5}_{31},\ldots,\widetilde{a}^{5}_{3\widetilde{r}_{5}},\widetilde{a}^{6}_{31},\ldots,\widetilde{a}^{6}_{3\widetilde{r}_{6}}]$, respectively. Then $h^{2}$ is a rational map with branching datum $[2\widetilde{a}^{3}_{31},\ldots,2\widetilde{a}^{3}_{3\widetilde{r}_{3}}\\ ,2\widetilde{a}^{4}_{31},\ldots,2\widetilde{a}^{4}_{3\widetilde{r}_{4}}]$, $[2\widetilde{a}^{5}_{31},\ldots,2\widetilde{a}^{5}_{3\widetilde{r}_{5}},2\widetilde{a}^{6}_{31},\ldots,2\widetilde{a}^{6}_{3\widetilde{r}_{6}}]$ and $[2\widetilde{a}^{1}_{31},\ldots,2\widetilde{a}^{1}_{3\widetilde{r}_{1}},2\widetilde{a}^{2}_{31},\\ \ldots,2\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$.
By Lemma 3.3, there there exists a rational map $\tau:S^{2}\rightarrow S^{2}$ with branching points at 00 and $\infty$ corresponding the partitions $[a_{11},\ldots,a_{1r_{1}}]$ and $[a_{21},\ldots,a_{2r_{2}}]$, branching points at $1,e^{\frac{2\pi}{3}\sqrt{-1}}$ and $e^{\frac{2\pi}{3}\sqrt{-1}}$ corresponding the partitions $[2\widetilde{a}^{3}_{31},\ldots,2\widetilde{a}^{3}_{3\widetilde{r}_{3}},2\widetilde{a}^{4}_{31},\ldots,2\widetilde{a}^{4}_{3\widetilde{r}_{4}}]$, $[2\widetilde{a}^{5}_{31},\ldots,2\widetilde{a}^{5}_{3\widetilde{r}_{5}},2\widetilde{a}^{6}_{31},\ldots,2\widetilde{a}^{6}_{3\widetilde{r}_{6}}]$  and $[2\widetilde{a}^{1}_{31},\ldots,2\widetilde{a}^{1}_{3\widetilde{r}_{1}},2\widetilde{a}^{2}_{31},\ldots,2\widetilde{a}^{2}_{3\widetilde{r}_{2}}]$, and other branching points assigned suitable complex numbers. By direct calculation, $\tau^{3}$ is the desired map. ∎
4. examples
As an application of our findings, we present some new exceptional examples when the number of branching points is at least three.
Example 4.1.
Suppose $d=2k,k\geq 4$, then the candidate branching datum of $(d,\{[k+2,\underbrace{1,\ldots,1}_{k-2}],[k-2,\underbrace{1,\ldots,1}_{k+2}],[\underbrace{2,\ldots,2}_{k}],[\underbrace{2,\ldots,2}_{k}]\})$ is exceptional.
Proof.
According to Corollary 1.1, the data is exceptional. ∎
Example 4.2.
Suppose $d=8k,k\geq 5$, then the candidate branching datum of $(d,\{[k+1,\underbrace{1,\ldots,1}_{7k-1}],[k-3,\underbrace{1,\ldots,1}_{7k+3}],[3,\underbrace{1,\ldots,1}_{8k-3}],[\underbrace{8,\ldots,8}_{k}],[\underbrace{8,\ldots,8}_{k}]\})$ is exceptional.
Proof.
Since $k+1>k$, according to Corollary 1.1 the data is exceptional. ∎
Example 4.3.
Suppose $d=10k,k\geq 9$, then the candidate branching datum of $(d,\{[k+1,\underbrace{1,\ldots,1}_{9k-1}],[k-7,\underbrace{1,\ldots,1}_{9k+7}],[3,\underbrace{1,\ldots,1}_{10k-3}],[3,\underbrace{1,\ldots,1}_{10k-3}],[3,\underbrace{1,\ldots,1}_{10k-3}]$, $[\underbrace{10,\ldots,10}_{k}],[\underbrace{10,\ldots,10}_{k}]\})$ is exceptional.
Proof.
Since $k+1>k$, according to Corollary 1.1 the data is exceptional. ∎
Example 4.4.
In the following tables, we give some new exceptional branching datum:
$n=4$ Some examples of exceptional branching datum
$d=2k,k\geq 4$ $\mathcal{D}=\{[k+2,\underbrace{1,\ldots,1}_{k-2}],[k-2,\underbrace{1,\ldots,1}_{k+2}],[\underbrace{2,\ldots,2}_{k}],[\underbrace{2,\ldots,2}_{k}]\}$
$d=3k,k\geq 3$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{2k-1}],[k-1,\underbrace{1,\ldots,1}_{2k+1}],[\underbrace{3,\ldots,3}_{k}],[\underbrace{3,\ldots,3}_{k}]\}$
$d=4k,k\geq 3$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{3k-1}],[k-1,\underbrace{1,\ldots,1}_{3k+1}],[\underbrace{4,\ldots,4}_{k}],[\underbrace{4,\ldots,4}_{k}]\}$
$n=5$ Some examples of exceptional branching datum
$d=6k,k\geq 4$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{5k-1}],[k-2,\underbrace{1,\ldots,1}_{5k+2}],[2,\underbrace{1,\ldots,1}_{6k-2}],[\underbrace{6,\ldots,6}_{k}],[\underbrace{6,\ldots,6}_{k}]\}$
$d=6k,k\geq 6$ $\mathcal{D}=\{[k+1,k-5,\underbrace{1,\ldots,1}_{4k+4}],[3,\underbrace{1,\ldots,1}_{6k-3}],[3,\underbrace{1,\ldots,1}_{6k-3}],[\underbrace{6,\ldots,6}_{k}],[\underbrace{6,\ldots,6}_{k}]\}$
$d=7k,k\geq 1$ $\mathcal{D}=\{[2k,\underbrace{1,\ldots,1}_{5k+5}],[3,\underbrace{1,\ldots,1}_{7k-3}],[3,\underbrace{1,\ldots,1}_{7k-3}],[\underbrace{7,\ldots,7}_{k}],[\underbrace{7,\ldots,7}_{k}]\}$
$d=7k,k\geq 7$ $\mathcal{D}=\{[k+2,k-6,\underbrace{1,\ldots,1}_{5k+4}],[3,\underbrace{1,\ldots,1}_{7k-3}],[3,\underbrace{1,\ldots,1}_{7k-3}],[\underbrace{7,\ldots,7}_{k}],[\underbrace{7,\ldots,7}_{k}]\}$
$d=8k,k\geq 9$ $\mathcal{D}=\{[k+2,k-8,\underbrace{1,\ldots,1}_{6k+6}],[4,\underbrace{1,\ldots,1}_{8k-4}],[4,\underbrace{1,\ldots,1}_{8k-4}],[\underbrace{8,\ldots,8}_{k}],[\underbrace{8,\ldots,8}_{k}]\}$
$d=8k,k\geq 5$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{7k-1}],[k-3,\underbrace{1,\ldots,1}_{7k+3}],[3,\underbrace{1,\ldots,1}_{8k-3}],[\underbrace{8,\ldots,8}_{k}],[\underbrace{8,\ldots,8}_{k}]\}$
$n=6$ Some examples of exceptional branching datum
$d=8k,k\geq 3$ $\mathcal{D}=\{[k+2,k-2,\underbrace{1,\ldots,1}_{6k+6}],[3,\underbrace{1,\ldots,1}_{8k-3}],[3,\underbrace{1,\ldots,1}_{8k-3}],[3,\underbrace{1,\ldots,1}_{8k-3}],[\underbrace{8,\ldots,8}_{k}],[\underbrace{8,\ldots,8}_{k}]\}$
$d=8k,k\geq 7$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{7k-1}],[k-5,\underbrace{1,\ldots,1}_{7k+5}],[3,\underbrace{1,\ldots,1}_{8k-3}],[3,\underbrace{1,\ldots,1}_{8k-3}],[\underbrace{8,\ldots,8}_{k}],[\underbrace{8,\ldots,8}_{k}]\}$
$d=9k,k\geq 8$ $\mathcal{D}=\{[k+2,\underbrace{1,\ldots,1}_{8k-2}],[k-6,\underbrace{1,\ldots,1}_{8k+6}],[3,\underbrace{1,\ldots,1}_{9k-3}],[3,\underbrace{1,\ldots,1}_{9k-3}],[\underbrace{9,\ldots,9}_{k}],[\underbrace{9,\ldots,9}_{k}]\}$
$d=9k,k\geq 8$ $\mathcal{D}=\{[k+1,k-7,\underbrace{1,\ldots,1}_{7k+6}],[3,\underbrace{1,\ldots,1}_{9k-3}],[3,\underbrace{1,\ldots,1}_{9k-3}],[3,\underbrace{1,\ldots,1}_{9k-3}],[\underbrace{9,\ldots,9}_{k}],[\underbrace{9,\ldots,9}_{k}]\}$
$d=10k,k\geq 7$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{9k-1}],[k-5,\underbrace{1,\ldots,1}_{9k+5}],[3,\underbrace{1,\ldots,1}_{10k-3}],[3,\underbrace{1,\ldots,1}_{10k-3}],[\underbrace{10,\ldots,10}_{k}],[\underbrace{10,\ldots,10}_{k}]\}$
$d=10k,k\geq 5$ $\mathcal{D}=\{[k+1,k-7,\underbrace{1,\ldots,1}_{8k+6}],[3,\underbrace{1,\ldots,1}_{10k-3}],[3,\underbrace{1,\ldots,1}_{10k-3}],[3,\underbrace{1,\ldots,1}_{10k-3}],[\underbrace{10,\ldots,10}_{k}],[\underbrace{10,\ldots,10}_{k}]\}$
$n=7$ Some examples of exceptional branching datum
$d=4k,k\geq 9$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{3k-1}],[k-7,\underbrace{1,\ldots,1}_{3k+7}],[3,\underbrace{1,\ldots,1}_{4k-3}],[3,\underbrace{1,\ldots,1}_{4k-3}],[3,\underbrace{1,\ldots,1}_{4k-3}],[\underbrace{4,\ldots,4}_{k}],[\underbrace{4,\ldots,4}_{k}]\}$
$d=4k,k\geq 10$ $\mathcal{D}=\{[k+1,k-9,\underbrace{1,\ldots,1}_{2k+8}],[3,\underbrace{1,\ldots,1}_{4k-3}],[3,\underbrace{1,\ldots,1}_{4k-3}],[3,\underbrace{1,\ldots,1}_{4k-3}],[3,\underbrace{1,\ldots,1}_{4k-3}],[\underbrace{4,\ldots,4}_{k}],[\underbrace{4,\ldots,4}_{k}]\}$
$d=5k,k\geq 12$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{4k-1}],[k-10,\underbrace{1,\ldots,1}_{4k+10}],[4,\underbrace{1,\ldots,1}_{5k-4}],[4,\underbrace{1,\ldots,1}_{5k-4}],[4,\underbrace{1,\ldots,1}_{5k-4}],[\underbrace{5,\ldots,5}_{k}],[\underbrace{5,\ldots,5}_{k}]\}$
$d=5k,k\geq 14$ $\mathcal{D}=\{[k+1,k-13,\underbrace{1,\ldots,1}_{3k+12}],[4,\underbrace{1,\ldots,1}_{5k-4}],[4,\underbrace{1,\ldots,1}_{5k-4}],[4,\underbrace{1,\ldots,1}_{5k-4}],[4,\underbrace{1,\ldots,1}_{5k-4}],[\underbrace{5,\ldots,5}_{k}],[\underbrace{5,\ldots,5}_{k}]\}$
$d=6k,k\geq 6$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{5k-1}],[k-4,\underbrace{1,\ldots,1}_{5k+4}],[2,\underbrace{1,\ldots,1}_{6k-2}],[2,\underbrace{1,\ldots,1}_{6k-2}],[2,\underbrace{1,\ldots,1}_{6k-2}],[\underbrace{6,\ldots,6}_{k}],[\underbrace{6,\ldots,6}_{k}]\}$
$d=6k,k\geq 6$ $\mathcal{D}=\{[k+1,k-5,\underbrace{1,\ldots,1}_{4k+4}],[2,\underbrace{1,\ldots,1}_{6k-2}],[2,\underbrace{1,\ldots,1}_{6k-2}],[2,\underbrace{1,\ldots,1}_{6k-2}],[2,\underbrace{1,\ldots,1}_{6k-2}],[\underbrace{6,\ldots,6}_{k}],[\underbrace{6,\ldots,6}_{k}]\}$
$d=7k,k\geq 6$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{6k-1}],[k-4,\underbrace{1,\ldots,1}_{6k+4}],[2,\underbrace{1,\ldots,1}_{7k-2}],[2,\underbrace{1,\ldots,1}_{7k-2}],[2,\underbrace{1,\ldots,1}_{7k-2}],[\underbrace{7,\ldots,7}_{k}],[\underbrace{7,\ldots,7}_{k}]\}$
$d=7k,k\geq 6$ $\mathcal{D}=\{[k+1,k-5,\underbrace{1,\ldots,1}_{5k+4}],[2,\underbrace{1,\ldots,1}_{7k-2}],[2,\underbrace{1,\ldots,1}_{7k-2}],[2,\underbrace{1,\ldots,1}_{7k-2}],[2,\underbrace{1,\ldots,1}_{7k-2}],[\underbrace{7,\ldots,7}_{k}],[\underbrace{7,\ldots,7}_{k}]\}$
$d=9k,k\geq 6$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{8k-1}],[k-4,\underbrace{1,\ldots,1}_{8k+4}],[2,\underbrace{1,\ldots,1}_{9k-2}],[2,\underbrace{1,\ldots,1}_{9k-2}],[2,\underbrace{1,\ldots,1}_{9k-2}],[\underbrace{9,\ldots,9}_{k}],[\underbrace{9,\ldots,9}_{k}]\}$
$d=9k,k\geq 6$ $\mathcal{D}=\{[k+1,k-5,\underbrace{1,\ldots,1}_{7k+4}],[2,\underbrace{1,\ldots,1}_{9k-2}],[2,\underbrace{1,\ldots,1}_{9k-2}],[2,\underbrace{1,\ldots,1}_{9k-2}],[2,\underbrace{1,\ldots,1}_{9k-2}],[\underbrace{9,\ldots,9}_{k}],[\underbrace{9,\ldots,9}_{k}]\}$
$d=10k,k\geq 9$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{9k-1}],[k-7,\underbrace{1,\ldots,1}_{9k+7}],[3,\underbrace{1,\ldots,1}_{10k-3}],[3,\underbrace{1,\ldots,1}_{10k-3}],[3,\underbrace{1,\ldots,1}_{10k-3}],[\underbrace{10,\ldots,10}_{k}],[\underbrace{10,\ldots,10}_{k}]\}$
$d=10k,k\geq 14$ $\mathcal{D}=\{[k+1,k-13,\underbrace{1,\ldots,1}_{8k+12}],[3,\underbrace{1,\ldots,1}_{10k-4}],[3,\underbrace{1,\ldots,1}_{10k-4}],[3,\underbrace{1,\ldots,1}_{10k-4}],[3,\underbrace{1,\ldots,1}_{10k-4}],[\underbrace{10,\ldots,10}_{k}],[\underbrace{10,\ldots,10}_{k}]\}$
$n=8$ Some examples of exceptional branching datum
$d=2k,k\geq 12$ $\mathcal{D}=\{[k+2,\underbrace{1,\ldots,1}_{k-2}],[k-10,\underbrace{1,\ldots,1}_{k+10}],[3,\underbrace{1,\ldots,1}_{2k-3}],[3,\underbrace{1,\ldots,1}_{2k-3}],[3,\underbrace{1,\ldots,1}_{2k-3}],[3,\underbrace{1,\ldots,1}_{2k-3}],[\underbrace{2,\ldots,2}_{k}],[\underbrace{2,\ldots,2}_{k}]\}$
$d=3k,k\geq 7$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{2k-1}],[k-5,\underbrace{1,\ldots,1}_{2k+5}],[2,\underbrace{1,\ldots,1}_{3k-2}],[2,\underbrace{1,\ldots,1}_{3k-2}],[2,\underbrace{1,\ldots,1}_{3k-2}],[2,\underbrace{1,\ldots,1}_{3k-2}],[\underbrace{3,\ldots,3}_{k}],[\underbrace{3,\ldots,3}_{k}]\}$
$d=4k,k\geq 7$ $\mathcal{D}=\{[k+1,\underbrace{1,\ldots,1}_{3k-1}],[k-5,\underbrace{1,\ldots,1}_{3k+5}],[2,\underbrace{1,\ldots,1}_{4k-2}],[2,\underbrace{1,\ldots,1}_{4k-2}],[2,\underbrace{1,\ldots,1}_{4k-2}],[2,\underbrace{1,\ldots,1}_{4k-2}],[\underbrace{4,\ldots,4}_{k}],[\underbrace{4,\ldots,4}_{k}]\}$
$d=4k,k\geq 7$ $\mathcal{D}=\{[k+1,k-6,\underbrace{1,\ldots,1}_{2k+5}],[2,\underbrace{1,\ldots,1}_{4k-2}],[2,\underbrace{1,\ldots,1}_{4k-2}],[2,\underbrace{1,\ldots,1}_{4k-2}],[2,\underbrace{1,\ldots,1}_{4k-2}],[2,\underbrace{1,\ldots,1}_{4k-2}],[\underbrace{4,\ldots,4}_{k}],[\underbrace{4,\ldots,4}_{k}]\}$
Acknowledgments
We are deeply grateful to Professor Zizhou Tang for his valuable insights and constructive advice, which greatly assisted in refining the quality and clarity of our initial submission.
Declarations
Data Availability Statement This manuscript has no associated data.
Competing interests On behalf of all authors, the corresponding author declares that there is no conflict of interest.